# Robust Training of Neural Networks at Arbitrary Precision and Sparsity

### v2

@@ 6.4 Scaling to Larger Models: A 4B Parameter Case Study
To investigate whether a larger model, when quantized to a similar footprint, can outperform a smaller model, we conducted a 26B token pre-training run comparing the 700M-parameter Gemma3-1B against the 3.2B-parameter Gemma3-4B.
@@ 6.4.1 Storage and Compute Efficiency at Scale
Our results, visualized in Figure 2, demonstrate that a larger model aggressively quantized with our method can be superior to a smaller model, even one at higher precision. The Gemma3-4B model, quantized to A4W1 with 2:4 sparsity, achieves higher accuracy (0.4517) than both the BF16 Gemma3-1B (0.4494) and the quantized A4W4 Gemma3-1B (0.4443). The energy efficiency analysis reveals an even clearer advantage: the sparse, quantized 4B model is not only more accurate than a quantized 1B model but achieves this with a significantly lower total computational cost.
Figure 2: (a) Storage vs. Accuracy comparison between Gemma3 1B and Gemma3 4B models. The quantized 4B model achieves higher accuracy than both BF16 and quantized versions of the 1B model. (b) Total Energy Cost vs. Accuracy. The quantized and sparse 4B model is both more accurate and more computationally efficient than a quantized 1B model.
@@ 6.4.2 Scaling Properties and Hardware Implications
These scaling experiments suggest a powerful principle: for a given storage or compute budget, training a larger, more redundant model and then applying aggressive quantization and sparsity is a more effective path to high performance than training a smaller model at higher precision. The true benefit extends beyond these metrics to hardware implementation. Ultra-low-precision models, particularly those operating at 1-bit or ternary level, can be executed on radically simpler hardware. Expensive floating-point units can be replaced with highly efficient bitwise operation circuits, leading to significant reductions in power consumption, silicon area, and cost. This opens a path for deploying state-of-the-art intelligence on edge devices and specialized accelerators. After all, biological systems achieve remarkable intelligence using discrete, spike-based (effectively binary) signals. Our findings suggest that as artificial neural networks continue to scale, they too can leverage the profound efficiency of low-precision signals to deliver superb performance.
@@ 7 Summary
For years, training hyper-efficient models has been hampered by the non-differentiable nature of quantization and sparsification. We argue the primary culprit is the Straight-Through Estimator’s (STE) quantization-oblivious backward pass—a critical blind spot that fuels training instability. We solve this with a denoising dequantization transform, born from a principled ridge regression objective, that forces the learning process to become robust to quantization error. This establishes a single, reliable framework for training at any precision or sparsity using off-the-shelf recipes, rendering empirical hacks obsolete. By making hyper-efficient, low-precision models a practical reality, our framework helps pave the way toward AI that mirrors the efficiency of biological intelligence and can be deployed ubiquitously.
@@ Appendix A Appendix
A.1 Related Work
Our work addresses the core problem of training instability from a different, foundational perspective. Instead of developing bit-specific heuristics or entirely new gradient estimators, we explicitly model the quantization discontinuity as an additive perturbation. Our primary contribution is a denoising dequantization transform, derived from a principled objective, that makes the training process inherently aware of and robust to this quantization error. This paper demonstrates that our approach provides a stable and effective foundation for training today’s LLMs at extreme low precision, often without the complex modifications proposed in related works.
A.2 Ablation Studies and Analysis
This appendix provides a comprehensive set of supplementary experiments and analyses that offer a deeper validation of our framework’s components and its broad applicability. We begin with a series of detailed ablation studies: we analyze the consistent superiority of our method over the standard STE (Sec. A.2.1), quantify the benefits of affine quantization unlocked by our approach (Sec. A.2.2), and explore the application of advanced techniques, including our centered Hadamard transform (Sec. A.2.3), low-precision FP4 formats (Sec. A.2.4), and the trade-offs in structured sparsity (Sec. A.2.5). We explain how we measure the energy consumption of LLMs in Sec. A.2.6. To demonstrate the versatility of our method beyond the LLM pre-training context of the main paper, we conclude by presenting results on established benchmarks, including ResNet-50 on ImageNet (Sec. A.2.7) and Transformers on WMT machine translation tasks (Sec. A.2.8).
A.2.1 Our Method Consistently Outperforms STE
@@
A.2.2 Affine Quantization Gains are Unlocked by Our Method
@@
A.2.3 Centered Hadamard Transform for Outlier Mitigation
@@
This analysis highlights two distinct philosophies for handling the outlier dilemma (noise vs. feature). The rotation-based approach blends outliers into other coordinates, distributing their magnitude. Our framework, in contrast, blends outliers into the statistical parameters (scale and offset) of the dequantization transform by having all samples participate in the ridge regression (Eq. 3). This is made robust by SCQ, which ensures the statistical blending is a local effect. While our centered transform improves the Hadamard implementation (Table 3, Figure 5) , our main experiments (Figure 1(c)) show that the localized statistical blending via moderate-block-size SCQ consistently outperforms the gains from the global coordinate-blending of the Hadamard transform, confirming it as a more direct and effective solution.
@@
A.2.4 Quantization to Low-Precision Floats (FP4)
@@
A.2.5 Trade-offs in Structured Sparsity and Quantization
@@
A.2.6 The Energy Efficiency Frontier: Methodology
@@
It is worth noting that some techniques, such as the Hadamard transform, introduce additional computational costs not captured by this MAC-centric metric. While the Hadamard transform has a complexity of $O(n\log n)$, for simplicity, we omit a detailed analysis of its energy contribution in this study. This metric nevertheless provides a unified and principled foundation to systematically map the trade-offs between total energy efficiency and model accuracy. By using this score, we can construct the energy efficiency Pareto frontier shown in Figures 1(d) 2(b), revealing the synergistic power of combining our robust training framework with asymmetric quantization and structured sparsity.
@@ A.2.8 Transformer on WMT
Fixed-Point (Boo & Sung, 2020) 28.48 26.94 Y Y Y 4
@@
A.2.9 Subchannel Block Size
@@
A.2.10 The Dual Role of the Denoising Transform and Regularization ($\lambda$)
@@ 0 NaN
A.2.11 Extra Results
We present the loss curve of training the GPT2-small model on the Openwebtext dataset (Fig. 6).
Figure 6: GPT-2 training Loss on Openwebtext.
A.3 Computational Flow for Quantized Matrix Multiplication
@@
A.4 Complexity Analysis of the Affine Shortcut
@@
A.5 Reference Code
@@ Appendix B Acknowledgements
We extend our sincere gratitude to Abhijit Ogale, Dinghua Li, Jeff Dean, Jian Li, Kyuyeuan Kim, Rasmus Larsen, Sameer Agarwal, Sanjiv Kumar, Tammo Splank for their insights and supports.

### v3

@@ 6.4 Scaling to Larger Models: A 4B Parameter Case Study
To investigate whether a larger model, when quantized to a similar footprint, can outperform a smaller model, we conducted a 26B token pre-training run comparing the 700M-parameter Gemma3 1B against the 3.2B-parameter Gemma3 4B.
@@ 6.4.1 Storage and Compute Efficiency at Scale
Our results, visualized in Figure 4, demonstrate that a larger model aggressively quantized with our method can be superior to a smaller model, even one at higher precision. The Gemma3 4B model, quantized to A4W1 with 2:4 sparsity, achieves higher accuracy (0.4517) than both the BF16 Gemma3 1B (0.4494) and the quantized A4W4 Gemma3 1B (0.4443). The energy efficiency analysis reveals an even clearer advantage: the sparse, quantized 4B model is not only more accurate than a quantized 1B model but achieves this with a significantly lower total computational cost. For a detailed tabular comparison across all 1B and 4B configurations, please refer to Table 6 in Appendix A.1.6.
Figure 4: (a) Storage vs. Accuracy comparison between Gemma3 1B and 4B models. The quantized 4B model achieves higher accuracy than both BF16 and quantized versions of the 1B model. (b) Total Energy Cost vs. Accuracy. The quantized and sparse 4B model is both more accurate and more computationally efficient than a quantized 1B model.
@@ 6.4.2 Scaling Properties and Hardware Implications
These experiments confirm that for a fixed budget, aggressively quantizing a larger model yields superior performance compared to a smaller, high-precision model. This trade-off offers distinct advantages for hardware implementation. Ultra-low-precision models, particularly those operating at the 1-bit or ternary level, can be executed on significantly simplified hardware architectures. By replacing complex floating-point units with efficient bitwise operation circuits, we can achieve substantial reductions in power consumption, silicon area, and manufacturing costs. This efficiency paves the way for deploying high-capacity models on edge devices and specialized accelerators where energy constraints are paramount.
@@ 7 Summary
The non-differentiable nature of quantization has long challenged efficient model training, as the community has viewed it as inherently unfriendly to gradient descent’s reliance on smoothness. We pinpoint—for the first time—that the core issue is the lack of a proper gradient path allowing training to learn robustness to quantization noise, which STE exacerbates by ignoring error in the backward pass. Our work provides a theoretically sound resolution through a ridge regression-derived denoising dequantization transform, ensuring well-defined gradients without heuristic estimation. This unified framework enables robust training at arbitrary precisions and sparsity with standard recipes and no ad-hoc fixes, unlocking the full potential of low-precision computing for resource-constrained applications.
@@ Appendix A Appendix
This appendix presents supplementary experiments and analyses that further validate our framework’s robustness and versatility. We begin with detailed ablation studies: we analyze our method’s consistent superiority over standard STE (Sec. A.1.1), quantify the gains unlocked by affine quantization (Sec. A.1.2), and explore advanced techniques including our centered Hadamard transform (Sec. A.1.3), low-precision FP4 formats (Sec. A.1.4), and structured sparsity trade-offs (Sec. A.1.5). Additionally, we outline our energy consumption methodology in Sec. A.1.6. To demonstrate applicability beyond LLM pre-training, we present results on ResNet-50 (Sec. A.1.7) and WMT machine translation tasks (Sec. A.1.8).
On the implementation side, we detail the computational flow of quantized matrix multiplication in Sec. A.2, analyze the complexity of the affine shortcut in Sec. A.3 and provide a rigorous gradient analysis in Sec. A.4. Finally, we provide an extended discussion of related work in Sec. A.5 and include reference code in Sec. A.6.
A.1 Ablation Studies and Analysis
A.1.1 Our Method Consistently Outperforms STE
@@
A.1.2 Affine Quantization Gains are Unlocked by Our Method
@@
A.1.3 Centered Hadamard Transform for Outlier Mitigation
@@
This analysis highlights two distinct philosophies for handling the outlier dilemma (noise vs. feature). The rotation-based approach blends outliers into other coordinates, distributing their magnitude. Our framework, in contrast, blends outliers into the statistical parameters (scale and offset) of the dequantization transform by having all samples participate in the ridge regression (Eq. 3). This is made robust by SCQ, which ensures the statistical blending is a local effect. While our centered transform improves the Hadamard implementation (Table 3, Figure 7) , our main experiments (Figure 2 show that the localized statistical blending via moderate-block-size SCQ consistently outperforms the gains from the global coordinate-blending of the Hadamard transform, confirming it as a more direct and effective solution.
@@
A.1.4 Quantization to Low-Precision Floats (FP4)
@@
A.1.5 Trade-offs in Structured Sparsity and Quantization
@@
A.1.6 The Energy Efficiency Frontier: Methodology
@@
It is worth noting that some techniques, such as the Hadamard transform, introduce additional computational costs not captured by this MAC-centric metric. While the Hadamard transform has a complexity of $O(n\log n)$, for simplicity, we omit a detailed analysis of its energy contribution in this study. This metric nevertheless provides a unified and principled foundation to systematically map the trade-offs between total energy efficiency and model accuracy. By using this score, we can construct the energy efficiency Pareto frontier shown in Figures 34, revealing the synergistic power of combining our robust training framework with asymmetric quantization and structured sparsity.
@@ A.1.8 Transformer on WMT
Fixed-Point (Boo and Sung, 2020) 28.48 26.94 Y Y Y 4
@@
A.1.9 Subchannel Block Size
@@
A.1.10 The Dual Role of the Denoising Transform and Regularization ($\lambda$)
@@ 0 NaN
A.1.11 Additional Experimental Validation
In this section, we provide further empirical evidence to substantiate the robustness of our method across different scales and distinct quantization schemes.
Scalability to OpenWebText (GPT-2 Small).
To validate that our stability findings generalize to standard language modeling benchmarks, we trained a GPT-2 Small model on the OpenWebText dataset. Figure 8 illustrates the training loss trajectories in the challenging A1W1 regime. Our method demonstrates smooth and stable convergence throughout the training process. In contrast, the standard STE baseline and BitNet exhibit significant instability or failure to converge to a competitive loss. This confirms that the robustness of our denoising estimator holds as model capacity and dataset complexity increase.
Figure 8: Scalability: GPT-2 Small on OpenWebText (A1W1). Our method (Blue) converges smoothly, whereas STE (Orange) and BitNet (Red Cross) exhibit instability or divergence.
Isolating Estimator Gains: Linear vs. Linear Comparison.
To rigorously isolate the contribution of our denoising estimator from the benefits of affine quantization (addressing Reviewer BVSw’s inquiry), we reproduced the Shakespeare A1W1 experiment with additional “linear-only” configurations. We compared our method against the standard STE, BitNet, and ParetoQ, restricting all methods to identical symmetric linear quantization schemes where applicable.
The results, presented in Figure 9, demonstrate three critical findings :
  1. 1.
    Estimator Stability (Linear vs. Linear): When restricted to linear quantization, our method converges stably to a low validation loss ($\approx 2.1$). In stark contrast, the standard STE with linear quantization exhibits significant instability and diverges to a high loss (>5.0>5.0). This confirms that our estimator provides a fundamental stability advantage independent of the quantization mapping.
  2. 2.
    Comparison with Strong Baselines: Our linear implementation outperforms both BitNet and ParetoQ in terms of convergence speed and final validation loss.
  3. 3.
    The Affine Advantage: While our linear method is robust, enabling our full affine quantization yields the lowest overall loss ($\approx 1.9$). This reinforces that our estimator is robust enough to unlock the additional expressivity of affine parameters, which other methods fail to utilize effectively.
Figure 9: Mechanism Isolation: Linear-vs-Linear Ablation (Shakespeare). We isolate the estimator’s effect by running “Ours” and “STE” with identical linear schemes. Ours (Linear) remains stable and outperforms STE (Linear), confirming the estimator’s intrinsic value. Ours (Affine) achieves the best overall performance.
A.2 Computational Flow for Quantized Matrix Multiplication
@@
A.3 Complexity Analysis of the Affine Shortcut
@@
A.4 Gradient Analysis: Explicit Dependence on Quantization Error
In Section 3.3.2, we posit that our denoising transform establishes a backward pass that is explicitly aware of the quantization error $\bm{\delta}$, addressing the ”blind spot” of the standard Straight-Through Estimator (STE). Here, we provide a rigorous derivation of the local gradient (Jacobian) to validate this claim.
Consider the symmetric (linear) dequantization case for simplicity (Eq. 4). The transform is defined as:
$g(\bm{q})=s_{g}(\bm{q})\cdot\bm{q}$ (10)
where the scale factor $s_{g}$ is not a constant parameter, but a scalar function of the quantized vector $\bm{q}$ derived from the ridge regression solution: $s_{g}(\bm{q})=\frac{\bm{x}^{\top}\bm{q}}{\bm{q}^{\top}\bm{q}+\lambda}$.
To analyze the backward pass, we compute the Jacobian matrix $\bm{J}=\frac{\partial g(\bm{q})}{\partial\bm{q}}$. Applying the product rule to $g(\bm{q})$ yields:
$\frac{\partial g(\bm{q})}{\partial\bm{q}}=s_{g}(\bm{q})\cdot\mathbf{I}+\bm{q}\cdot\left(\nabla_{\bm{q}}s_{g}(\bm{q})\right)^{\top}$ (11)
where $\mathbf{I}$ is the identity matrix.
The first term, $s_{g}(\bm{q})\cdot\mathbf{I}$, corresponds to a standard scaling operation. The second term, however, introduces a data-dependent rank-1 correction. We derive $\nabla_{\bm{q}}s_{g}(\bm{q})$ using the quotient rule:
$\nabla_{\bm{q}}s_{g}=\frac{(\bm{q}^{\top}\bm{q}+\lambda)\bm{x}-(\bm{x}^{\top}\bm{q})2\bm{q}}{(\bm{q}^{\top}\bm{q}+\lambda)^{2}}$ (12)
Substituting this result back into the Jacobian expression reveals that the gradient is a complex, non-linear function of $\bm{q}$.
The Role of Quantization Error.
Since $\bm{q}=f(\bm{x})+\bm{\delta}$, the local derivative $\frac{\partial g}{\partial\bm{q}}$ is explicitly parameterized by the quantization error $\bm{\delta}$.
Comparison with STE.
In standard STE, the backward pass typically assumes an identity (or constant scalar) derivative $\frac{\partial g(\bm{q})}{\partial\bm{x}}\approx\mathbf{I}$, effectively rendering the gradient independent of the specific value of $\bm{\delta}$. In contrast, our method computes the gradient of the loss with respect to $\bm{x}$ via the chain rule:
$\frac{\partial\mathcal{L}}{\partial\bm{x}}=\frac{\partial\mathcal{L}}{\partial g(\bm{q})}\cdot\underbrace{\frac{\partial g(\bm{q})}{\partial\bm{q}}}_{\text{Error-Aware}}\cdot\frac{\partial\bm{q}}{\partial\bm{x}}$ (13)
Because the term $\frac{\partial g(\bm{q})}{\partial\bm{q}}$ varies based on the magnitude and direction of the noise $\bm{\delta}$ (embedded within $\bm{q}$), the learning signal flowing back to the weights is dynamically modulated by the instantaneous quantization error. This mechanism is mathematically analogous to the stabilizing gradients found in Normalization layers, forcing the network to adapt to the noisy quantization surface.
A.5 Related Work
Our work addresses the core problem of training instability from a different, foundational perspective. Instead of developing bit-specific heuristics or entirely new gradient estimators, we explicitly model the quantization discontinuity as an additive perturbation. Our primary contribution is a denoising dequantization transform, derived from a principled objective, that makes the training process inherently aware of and robust to this quantization error. By deriving well-defined gradients without surrogates, our approach offers a theoretically grounded alternative to the heuristic, gradient estimation-based methods.
A.6 Reference Code
@@ Appendix B Acknowledgements
We extend our sincere gratitude to Abhijit Ogale, Dinghua Li, Jeff Dean, Jian Li, Kyuyeuan Kim, Rasmus Larsen, Sameer Agarwal, Sanjiv Kumar, Sergey Ioffe, Tammo Splank, Zhifeng Chen for their insights and supports.